Suppose, the average U. S. household spends $90 per day. You recently took a random sample of 30 households in Huntsville and the results revealed a mean of $84.50. Suppose the SD is known to be $14.50. Using a 0.05 level of significance, can it be concluded that the average amount spent per day by U.S. households has decreased


1
Expert's answer
2021-05-31T16:37:22-0400

"\bar{x}=84.50 \\

\sigma=14.50 \\

n=30 \\

H_0: \mu= 90 \\

H_1: \mu<90"

The test statistic:

"Z=\frac{\bar{x}-\mu}{\sigma / \sqrt{n}} \\

Z = \frac{84.50-90}{14.50/ \sqrt{30}}=-2.08"

The level of significance α=0.05

Since the test is left tailed. Thus, the critical value using the standard normal table is -1.645.

Reject H0 if "Z<Z_{crit}" .

"Z=-2.08<Z_{crit}=-1.645"

Thus, reject the H0.

Hence, it can be concluded that the average amount spent per day by U.S. households has decreased.